OpenStudy (anonymous):

how many significant digit of 0.03

OpenStudy (noseboy908):

The was I was taught this was as follows: Think of how you could re-write the problem. Aka, think of how you could simplify it, even if it's just removing one unnecessary number. In this case, the first zero is insignificant because the number remains the same even if rewritten as ".03". Consider this though, can we remove the second zero, and keep the number as is?

OpenStudy (anonymous):

i am sure it will be one but my teacher said 3! since we can write 3*10^-2

OpenStudy (noseboy908):

Your teacher said the answer was three?

OpenStudy (anonymous):

he said has three significant digit ,but i think it is just one

OpenStudy (noseboy908):

I'm not sure where he got that answer, but I'd reason that it's 2, but here's why: you can remove the first zero, and keep the same value present, but not the second zero, and the three is always significant, so I'd suggest two as your answer.

OpenStudy (anonymous):

EXAMPLES 5.1 has 1 significant digit of 5: |5 − 5.1| = 0.1 0.51 has 1, not 2, significant digits of 0.5: |0.5 − 0.51| = 0.01 4.995 has 3 significant digits of 5: 5 − 4.995 = 0.005 4.994 has 2, not 3, significant digits of 5: 5 − 4.994 = 0.006 0.5 has all significant digits of 0.5 1.4 has 0 significant digits of 2: 2 − 1.4 = 0.6

OpenStudy (anonymous):

relative error Note that the tail portion of the form 5000 . . . = 4999 . . . is still allowed.